Solve for y, then find the value of y for the given value of x
3x-6y=6; x=2

Answer by checkley77(12569)

3x-6y=6
-6y=-3x+6
y=-3x/-6+6/-6
y=x/2-1
x=2 then:
y=2/20-1
y=1-1
y=0